# Balancing Class Accuracies for Language Models in Inference Time via Nonlinear Integer Programming

### v3

@@ 3.2 Solving DNIP with Simulated Annealing
We solve DNIP by simulated annealing [6, 7]. Simulated annealing (SA) is a Metropolis-Hastings sampling algorithm, which is widely adopted to solve optimization problems with its versatility in dealing with highly nonlinear models, noisy data, and many constraints [8]. The DNIP mathematical model is such a case, pointing to the need of SA. We especially take advantages of SA algorithms’ searching ability - escaping from local optima and finding global optima after reaching local optima - to find optimal correction weights.
In details, the SA strategy consists of searching the solution space starting from a randomly initialized $\xi$ and then generating a new one by perturbing it. The new solution is either accepted or rejected by the Metropolis criterion after evaluating the objectives.
Algorithm 1 describes the searching for optimal $\xi$ using SA. The outer loop performs the cooling/annealing process. In this work, a geometric decay with $\alpha=0.95$ and initial temperature of 200,000 is set as cooling schedule. The inner loop simulates the thermal equilibrium reaching process at a temperature, and its criterion ensures that either the number of generated or accepted solutions at a temperature is above a threshold [9]. We sample a new $\xi$ from the neighborhood by randomly substitute a selection in $\xi$. The computational complexity for the SA algorithm is $\mathcal{O}(NK)$. Please refer to Appendix C for derivations.
@@ 4.1 Experimental Setup
Evaluation Tasks. The proposed method is evaluated on a diverse set of binary and multi-class NLP classification tasks, across general domains and biomedical domains. The five general-domain evaluation datasets are 4-way news topic classification AGNews [10], 14-way ontology classification DBpedia [11], 5-way sentiment classification SST-5 [12], 6-way retrieval question classification TREC [13, 5], binary entailment recognition RTE [14]; the two domain-specific datasets from BLURB [15] are 5-way biomedical relation extraction DDI [16], and 3-way biomedical question answering PubMedQA [17]. The selected biomedical tasks are of practical uses. For example, detecting DDI is useful in preventing adverse effects from drug combinations.
Each dataset comes with a training and a test set. We make use of the default training set for DNIP optimization, except for AGNews, DBpedia, and DDI, for which we randomly select 10,000 training examples since the datasets are quite large. We then split the training examples into 95%, 5% optimization and development sets. We evaluate on the default test set except for AGNews, for which we evaluate on 5,000 randomly selected test examples. Evaluation metrics are COBias and accuracy.
Models and Experimental Configuration. We evaluate three open-source LLMs across varied scales and model families, including GPT-2-XL (1.5B parameters), Llama-2-7B (7B parameters), and Llama-2-13B (13B parameters). For main results, we use 1-shot prompting and obtain the ICL output class softmax probabilities. To standardize calculations, we normalize the softmax probabilities over all classes. For DNIP, we tune parameters of $\beta,\tau,K,u$ on the development set. To account for variance caused by different demonstrations, for each model and dataset, we perform three runs of prompting with different 1-shot demonstrations and obtain three sets of initial probabilities. For both ICL and DNIP, we report the average test accuracy and COBias over the three runs. The demonstration is randomly selected from training examples. All prompting is done on an NVIDIA A100 GPU. Simulated annealing executes on CPU (execution time depends on the optimization set size, 3 to 30 minutes on AMD EPYC 7742 CPU or slightly longer on local laptop CPU).
@@ 4.2 Main Results
The average test accuracy (in black) and COBias (in blue) with standard deviation over three runs are shown in Table 3. In all three LLMs, test COBias by ICL is large, with GPT-2-XL exhibiting the largest COBias (50.3%), showing that class accuracy imbalance exists in LLMs of different sizes and families. With DNIP, test COBias is significantly reduced, by an absolute 29%, 23%, and 26% for the three models respectively; meanwhile, there is an 16%, 12%, and 10% absolute improvement in test accuracy respectively. It is worth noting that for tasks whose ICL test accuracy is already in higher 80s, e.g., DBpedia, DNIP can reduce COBias while further improving the accuracy to 90s. In addition, we observe a very large ICL COBias at 82.4% in GPT-2-XL for RTE, reflecting severe over-predictions in class True. After applying DNIP, the COBias greatly reduces to 7.1%. That is, DNIP obtains over 75% absolute reduction in COBias for RTE, demonstrating DNIP’s capability for balancing very imbalanced class accuracies without hurting the overall accuracy much. Moreover, on the two biomedical tasks, DNIP brings an average absolute reduction of 31% and 13% in COBias, and an average absolute increase of 34% and 11% in accuracy, showing its effectiveness for domain-specific tasks.
Moreover, the 13B Llama model can suffer even higher COBias than its 7B counterpart on biomedical tasks, further showing that class accuracy imbalance does not go away as model sizes increase, and DNIP is effective to balance class accuracies for LLMs of all sizes.
Back to the AGNews example in Table 2, we present the DNIP debiased results in Table 2. Over-predictions in Business are greatly reduced, and the accuracy of Tech is significantly improved from 19% to 84%. These results show that DNIP strongly mitigates COBias and greatly improves the accuracy of lower-performing classes, while boosting overall accuracy.
Computational Time. Regarding computational time needed to perform DNIP optimization, on all datasets, the optimization is finished in the scale of minutes. Though computational time increases with more classes and a more fine-grained weight scale, it should be noted that the time elapsed in optimization is dozens of minutes even for tasks of more than 10 classes, and a weight scale of 30 is good enough for most of the datasets experimented in this paper. This demonstrates the practicality of DNIP in the computational time needed for performing optimization.
@@
k-shot (1 demon. per class) DNIP COBias $7.3_{0.5}$ $4.3_{0.7}$ $2.8_{1.5}$ $12.1_{5.5}$ $5.0_{3.3}$ $12.5_{4.3}$ $8.7_{1.2}$ 7.5
@@
4.3 More ICL Settings
We prompted Llama-2-13B in two additional settings: 5-shot, where each demonstration is randomly selected as in the 1-shot case; k-shot, where k is the number of classes and each class is represented by a demonstration example in the prompt. Table 4 presents results on seven benchmark datasets in both settings. For example, the relative average COBias reduction (comparing DNIP to ICL) is 65% and 74% for 1-shot and 5-shot cases respectively, and the relative average accuracy increase is 16% and 11% for 1-shot and 5-shot cases respectively, demonstrating DNIP’s capabilities with more shots. In addition, prompt design does contribute to different output probabilities, which could be helpful to provide different starting points for DNIP, so DNIP could optimize for better solutions. However, prompt engineering alone may not be most effective to solve the COBias issue. For example, compared to 1-shot prompting (Table 3), increasing the number of shots does not significantly further reduce COBias or boost accuracy; using more diverse demonstrations (k-shot) may help with COBias, but does not gain much higher accuracy than 1-shot. Therefore, using more shots does not help reduce COBias much, and DNIP significantly reduces COBias and improves accuracy in both 5-shot and k-shot settings, further showcasing its effectiveness.
4.4 DNIP Significantly Reduces Over-predictions, and Balances Class Accuracies across Varied Classification Tasks, as many as 14 Classes
Figure 2: Pairwise class accuracy differences on test sets before and after applying DNIP; the pinker the higher class accuracy difference, the greener the lower; average accuracy over three runs are used. Heatmaps after DNIP show nearly 0 accuracy differences for many class pairs.
@@ 4.3 More ICL Settings
We visualize pairwise class accuracy differences before and after applying DNIP to 1-shot, Llama-2-13B ICL results in Figure 2. In each heatmap, row $i$, column $j$ cell represents the absolute difference between the test accuracies of two classes in row $i,j$ , i.e., $|A_{i}-A_{j}|$. DNIP brings the accuracy differences down to close to zero for AGNews, DDI, DBpedia, and RTE. The lighter green heatmaps in the bottom row show that DNIP consistently reduces COBias across different NLP classification tasks.
Relieving Odd Classes for 14-Way Ontology Classification. We further investigate the 14-way classification task DBpedia. Before debiasing, the Nat. class exhibits the largest class accuracy difference to other classes. In fact, Nat. is the most under-predicted class, whose odd class is Trans. and second odd class is Bldg., as shown by the confusion matrix in Figure 3 (based on seed 0, 1-shot ICL run). DNIP greatly reduces false negatives for Nat., i.e., reduces wrong Trans. and Bldg. predictions for Nat.. Moreover, on an average over all three runs, the test accuracy for Nat. improves from 44% to 89%, and that for Trans. also increases from 78% to 87%, demonstrating that DNIP is effective in reducing class accuracy differences for tasks of as many as 14 classes. Furthermore, we present more heatmaps for smaller LLMs including Llama-2-7B and GPT-2-XL in Appendix D.
@@ 4.5 Optimizing COBiassingle
Using COBiassingle (Eq. 4) for DNIP optimization will focus on odd classes only. Figure 4 shows ablation results when replacing the COBias term in the DNIP objective with COBiassingle.
$\min\quad Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta\frac{1}{N}\displaystyle\sum\nolimits_{i}\bigg{|}A_{\textrm{odd}_{i}}-A_{c_{i}}\bigg{|}-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (14)
@@
Blue dots represent the respective evaluation metric score of the original 1-shot ICL Llama-2-13B outputs of the seven benchmark datasets; green dots are debiased results optimized by the COBiassingle DNIP objective (Eq. 14); red dots are debiased results optimized by the COBias DNIP objective.
Comparing DNIP with COBiassingle and DNIP with COBias, their performances are close, but DNIP with COBias achieves higher overall test accuracy and slightly better bias mitigation performances. In other words, DNIP with COBiassingle can significantly minimize accuracy differences between odd classes and ground-truth classes, but DNIP with COBias is even more effective when it comes to balancing all pairs of classes, not limited to the odd classes. In summary, optimizing either COBiassingle or COBias is an effective inference-time measure to reduce over-predictions in raw ICL outputs and boost accuracy for under-predicted classes.
4.6 Accuracy-related Objectives vs. the COBias Objective
We show that the integration of accuracy-related objectives and the COBias objective are indispensable in achieving the best of both worlds in Figure 5. DNIP objective function $z$ consists of three terms: the error rate term $z_{1}=\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}$; the COBias term $z_{2}={N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}|A_{i}-A_{j}|$; and the PMI term $z_{3}=\sum_{j=1}^{N}\textrm{PMI}_{j}$; the accuracy-related terms are $z_{1}$ and $z_{3}$. We ablate all ablation objective functions to perform DNIP: $z_{1},z_{2},z_{3},z_{1}+\beta z_{2},z_{1}-\tau z_{3},\beta z_{2}-\tau z_{3},z_{1}+\beta z_{2}-\tau z_{3}$. Ablations on GPT-2-XL and Llama-2-7B show similar findings (Appendix E). To make fair comparisons, we fix $\beta=2.7,\tau=0.2$ and weight scale $K=30$, which may not result in the same scores reported in Table 3. Among all ablations, $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias. For the criteria, lower COBias is our top priority; when COBias scores are similarly low, we prefer the one that achieves higher accuracy. Compared to $z_{1}+\beta z_{2}-\tau z_{3}$$z_{1}$: (1) Using $z_{1}$ or $z_{3}$ solely emphasizes improving accuracy; though obtaining higher or similar accuracy, they suffer a higher COBias. (2) Using $z_{1}-\tau z_{3}$ also obtains a higher/similar accuracy but a higher COBias. (3) Using $z_{2}$ solely focuses on mitigating COBias and is straightforwardly effective in reducing COBias. However, the shortcoming is obvious: a lower accuracy. (4) Using $z_{1}+\beta z_{2}$ or $\beta z_{2}-\tau z_{3}$ is promising with a similar/lower accuracy and a slightly lower COBias. We note that accuracy-related objectives alone increase accuracy by over 16% ($75\%-59\%$), which is greater than using the objective combining COBias, suggesting DNIP is effective for those who emphasize accuracy improvement.
4.7 Performance Improvements vs. Number of Optimization Examples
Figure 6: Relative test COBias reduction; higher is better.
Figure 7: Relative test accuracy increase; higher is better.
Moreover, we observe for RTE a declining trend of accuracy increase, which matches its growing trend of COBias reduction, reflecting the trade-off between the two objectives. We also note a slight decrease in accuracy improvement for SST-5 and DDI at full size, which matches their sudden increase in COBias reduction, verifying that DNIP has emergent COBias reduction abilities with thousands of optimization examples. Overall, DNIP improves both accuracy and COBias at 10 examples, and achieves acceptable COBias reduction and accuracy improvement at 100 examples, with test COBias reduced by 57% and test accuracy improved. For detailed scores, please refer to Appendix F.
@@ 5 Related Work
LLMs may fail in various ways. Analysis towards understanding LLM failure modes has gained remarkable insights on how they fail and how to identify/avoid failures. For example, understanding hallucinations in LLMs facilitates fact-checking evaluation metrics for LLMs.[18, 19, 20]. Jones and Steinhardt [21] reveals failure modes of LLMs which resemble cognitive biases and measures the failures on GPT-3.5 and Codex. Lu et al. [22] identify ICL order sensitivity that result in LLM prompting failures. Turpin et al. [23] shows that LLM reasoning abilities are affected by adding biasing features to the input. Along this line, we tackle a common LLM failure of the class accuracy difference, via nonlinear integer programming optimization on the evaluation metric scores. The inference-time class accuracy bias has multi-faceted causes. It could also be attributed to pre-training data, such as predilections for common tokens in the pretraining corpora [24]. It could be a direct outcome of LLMs’ liking for specific patterns or answers, leading to highly imbalanced predicted class distributions [25]. Other causes include choices of prompt templates [26, 27, 28], demonstrations [29], and prompt orders [22, 23]. Unintended LLM behaviors may also exacerbate the bias, such as user-following sycophancy [30, 31, 32] and prompt toning effects [21, 33, 34, 35].

### v5

@@ 3.2 Solving DNIP with Simulated Annealing
Because the NIP objective contains discontinuous functions and is non-differentiable, we solve DNIP by simulated annealing [4, 5]. Simulated annealing (SA) is a Metropolis-Hastings sampling algorithm, which is widely adopted to solve optimization problems with its versatility in dealing with nonlinear models and many constraints [6]. The DNIP mathematical model is nonlinear, so SA is suitable. The main advantage of SA is escaping from local optima and finding global optima after reaching local optima, which is helpful in finding optimal correction weights.
In details, the SA algorithm consists of searching the solution space starting from a randomly initialized $\xi$ and generating a new one by perturbing it. The new solution is either accepted or rejected by the Metropolis criterion after evaluating the DNIP objective. Algorithm 1 describes the searching for optimal $\xi$ using SA. The outer loop performs the cooling/annealing process. In this work, the cooling schedule is set with a geometric decay with $\alpha=0.95$ and initial temperature of 200,000. The inner loop simulates the thermal equilibrium reaching process at a temperature, and the criterion ensures that either the number of generated or accepted solutions at a temperature is above a threshold [7]. We sample a new $\xi$ from the neighborhood by randomly substitute a selection in $\xi$. The computational complexity for SA is $\mathcal{O}(NK)$. Please refer to Appendix C for derivations.
@@ 4.1 Experimental Setup
Evaluation Tasks. We evaluate on a diverse set of binary and multi-class NLP classification tasks, across general domains and biomedical domains. The five general-domain evaluation datasets are 4-way news topic classification AGNews, 14-way ontology classification DBpedia [8], 5-way sentiment classification SST-5 [9], 6-way retrieval question classification TREC, binary entailment recognition RTE [10]; the two domain-specific datasets from BLURB [11] are 5-way biomedical relation extraction DDI [12], and 3-way biomedical question answering PubMedQA [13]. The selected biomedical tasks are of practical uses. For example, detecting DDI is useful in preventing adverse effects from drug combinations. Each dataset comes with a training and a test set. We take the full or subset of the training set and obtain ICL class probabilities with ground-truth labels as the optimization set for DNIP optimization. Evaluations are performed on each test set against COBias and overall accuracy. Optimization and dev splits can be found in Appendix D.
Models and Experimental Configuration. We evaluate three open-source LLMs across varied scales and families, including GPT-2-XL (1.5B parameters), Llama-2-7B, and Llama-2-13B. We do few-shot prompting on an NVIDIA A100 GPU, main results use 1-shot prompting, and we also perform more few-shot analyses. For DNIP, $\beta,\tau,K,\mu$ are tuned on the development set. To account for variance caused by demonstrations in prompts, we perform three runs for each dataset with different 1-shot demonstrations (randomly selected from the optimization set). Average scores over three runs are reported. SA executes on CPU.
@@ 4.2 Main Results
Averaged test accuracy (in black) and COBias (in blue) with standard deviation over three runs are shown in Table 3. Seen from the ICL COBias scores, class accuracy imbalance exists in LLMs of different sizes and families. With DNIP, test COBias is significantly reduced by an absolute 29%, 23%, and 26% for the three models respectively; test accuracy is improved by 16%, 12%, and 10% respectively. For tasks whose ICL test accuracy is already in higher 80s, e.g., DBpedia, DNIP can reduce COBias while further improving the accuracy to 90s. In addition, ICL COBias is very large (82.4%) on GPT-2-XL for RTE, reflecting severe under-predictions in class False. After applying corrections, its COBias greatly reduces to 7.1%, demonstrating DNIP’s capability for balancing very imbalanced class accuracies without hurting the overall accuracy much. DNIP is also effective for biomedical tasks, bringing an average absolute reduction of 31% and 13% in COBias, and an average absolute increase of 34% and 11% in accuracy.
Dataset Time/s Variable Size
AGNews 2278 120
DBpedia 2584 420
SST-5 1499 150
TREC 1274 180
RTE 518 60
DDI 2092 150
PubMedQA 203 90
Table 4: Optimization time (seconds).
Moreover, the 13B Llama model can suffer even higher COBias than its 7B counterpart on biomedical tasks, further showing that class accuracy imbalance does not go away as model sizes increase, and DNIP is effective to balance class accuracies for LLMs of all sizes. Furthermore, we present the confusion matrix of DNIP debiased AGNews results in Table 2. Under-predictions in Tech are greatly reduced, with its accuracy significantly improved from 19% to 84%. Additionally, DNIP requires only a small optimization set of 10 examples for a fair improvement in both COBias and accuracy, and it exhibits a further, emergent COBias reduction at several thousands optimization examples.
Figure 2: Objective function and temperature curves.
We discuss performance improvements vs. the number of optimization examples needed in Appendix E.
Computational Time. For all datasets, optimization finishes in minutes. Though the optimization time varies across variable size ($N\cdot K$) and optimization set size, it finishes in dozens of minutes even for DBpedia (14 classes; Figure 2). A weight scale of 30 is good enough for most datasets evaluated.
Optimization time using weight scale 30, $\beta=2.7,\tau=0.2$ is shown in Table 4, demonstrating the practicality of DNIP as an inference-time debiasing method without incurring much costs.
4.3 More ICL Settings
@@
DNIP COBias $7.3_{0.5}$ $4.3_{0.7}$ $2.8_{1.5}$ $12.1_{5.5}$ $5.0_{3.3}$ $12.5_{4.3}$ $8.7_{1.2}$ 7.5
@@
We prompt Llama-2-13B in two additional settings: 5-shot, where each demonstration is randomly selected; k-shot, where k is the number of classes and each class is represented by a demonstration, randomly selected. Table 5 shows that, for example, the relative average COBias reduction by DNIP is 65% and 74% for 1-shot (Table 3) and 5-shot respectively, with relative average accuracy increase of 16% and 11%. The rationale is, demonstrations contribute to variance in outputs, but prompt engineering alone does not mitigate much the imbalance. DNIP can correct the outputs no matter what different starting points are given.
@@
4.4 DNIP Significantly Balances Class Accuracies for Classification Tasks with as Many as 14 Classes
We further investigate DBpedia. By ICL, class Nat. exhibits the largest accuracy differences to others, and it is the most under-predicted class, whose odd class is Trans. and second odd class is Bldg., shown in Figure 3 (using seed 0 run). DNIP greatly reduces Trans. and Bldg. predictions for Nat.. When averaged over all three runs, test accuracy for Nat. improves from 44% to 89%, and that for Trans. increases from 78% to 87%, demonstrating DNIP’s competence in reducing class accuracy differences.
Figure 4: Pairwise class accuracy differences on test sets before and after DNIP; the pinker the higher difference, the greener the lower; averaged over three runs. DNIP achieves nearly 0 accuracy differences for many class pairs.
We also visualize pairwise class accuracy differences before and after applying DNIP to 1-shot, Llama-2-13B ICL results in Figure 4. In each heatmap, row $i$, column $j$ cell represents the absolute difference between the test accuracies of two classes in row $i,j$ , i.e., $|A_{c_{i}}-A_{c_{j}}|$. DNIP brings the accuracy differences down to close to zero for AGNews, DDI, DBpedia, and RTE. The lighter green heatmaps in the bottom row show that DNIP consistently reduces COBias across different NLP classification tasks. Furthermore, we present more heatmaps for smaller LLMs including Llama-2-7B and GPT-2-XL in Appendix F.
4.5 Accuracy-related Objectives vs. COBias Objective
We show that the integration of accuracy-related objectives and the COBias objective are indispensable in achieving the best of both worlds in Figure 5. DNIP objective function $z$ consists of three terms: the error rate term $z_{1}=\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}$; the COBias term $z_{2}={N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}|A_{c_{i}}-A_{c_{j}}|$; and the PMI term $z_{3}=\sum_{j=1}^{N}\textrm{PMI}_{j}$; the accuracy-related terms are $z_{1}$ and $z_{3}$. We ablate all ablation objective functions to perform DNIP: $z_{1},z_{2},z_{3},z_{1}+\beta z_{2},z_{1}-\tau z_{3},\beta z_{2}-\tau z_{3},z_{1}+\beta z_{2}-\tau z_{3}$. Ablations on GPT-2-XL and Llama-2-7B show similar findings (Appendix G). To make fair comparisons, we fix $\beta=2.7,\tau=0.2$ and weight scale $K=30$, which may not result in the same scores reported in Table 3. Among all ablations, $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias. For the criteria, lower COBias is our top priority; when COBias scores are similarly low, we prefer the one that achieves higher accuracy. Compared to $z_{1}+\beta z_{2}-\tau z_{3}$$z_{1}$: (1) Using $z_{1}$ or $z_{3}$ solely emphasizes improving accuracy; though obtaining higher or similar accuracy, they suffer a higher COBias. (2) Using $z_{1}-\tau z_{3}$ also obtains a higher/similar accuracy but a higher COBias. (3) Using $z_{2}$ solely focuses on mitigating COBias and is straightforwardly effective in reducing COBias, but the shortcoming is a lower accuracy. (4) Using $z_{1}+\beta z_{2}$ or $\beta z_{2}-\tau z_{3}$ is promising with a similar/lower accuracy and a lower COBias. Accuracy-related objectives alone increase accuracy by over 16% ($75\%-59\%$), suggesting DNIP is effective for those who emphasize accuracy improvement.
Figure 6: Comparisons to more methods.
@@ 4.6 Optimizing COBiassingle
Using COBiassingle (Eq. 3) for DNIP optimization will focus on odd classes only. Figure 7 shows ablation results when replacing the COBias term in the DNIP objective with COBiassingle. Blue dots represent the respective evaluation metric score of the original 1-shot ICL Llama-2-13B outputs of the seven benchmark datasets; green dots are debiased results optimized by the COBiassingle DNIP objective; red dots are results optimized by the COBias DNIP objective.
@@
Comparing DNIP with COBiassingle and DNIP with COBias, their performances are close, but DNIP with COBias achieves higher overall test accuracy and slightly better bias mitigation performances. In other words, DNIP with COBiassingle can significantly minimize accuracy differences between odd classes and ground-truth classes, but DNIP with COBias is even more effective when it comes to balancing all pairs of classes, not limited to the odd classes.
@@ 5 Related Work
Class accuracy imbalance is one of the potential failures of LLMs. Analysis towards understanding LLM failure modes has gained remarkable insights. For example, understanding hallucinations in LLMs facilitates fact-checking metrics for LLMs.[14, 15, 16]. Jones and Steinhardt [17] reveals failure modes of LLMs which resemble cognitive biases. Lu et al. [18] identify order sensitivity in prompting. The class accuracy imbalance has multi-faceted causes, from pre-training data, such as predilections for common tokens in pretraining [19], to a liking for specific patterns or answers [20]. Other causes include choices of prompt templates [21, 22, 23], demonstrations [24], and prompt orders [25]. Unintended LLM behaviors may also exacerbate it, such as sycophancy[26, 27, 28].